# Second- and third-order properties of multidimensional Langevin equations

### v13

@@ 2.4. Estimation of $\mathbf{A}$ from trajectories
We now consider the estimation of $\mathbf{A}$ in multiple dimensions. It has been shown that for estimation of the force field from stochastic trajectories, there is a finite rate at which information can be extracted [9]. This work has shown that the mean coefficient of the inferred force field $\widehat{\mathbf{A}}\mathbf{x}$ is $\langle\widehat{\mathbf{A}}\rangle=\mathbf{A}$ with correction $\mathcal{O}(T^{-1})$. However, they do not discuss the dependence of $\langle\widehat{\mathbf{A}}\rangle$ on dimension $d$. Eq. (C11) in [9] is summed over basis function index $\alpha$, and we may guess Eq. (C13) in [9] to be of order $\mathcal{O}(d^{3})$, leaving a $\mathcal{O}(d)$ bias for the inferred force field coefficient. We explicitly calculate the bias and variance of $\widehat{\mathbf{A}}$ for a $d$-dimensional linear system in which $-\mathbf{A}=\mathbf{D}=\mathbf{C}=\mathds{1}$. According to the information-theoretic criterion in [9], the force field should start to be resolved for a trajectory length of 2 time-units. The value of $\widehat{\mathbf{A}}$ inferred from a trajectory of duration $T$ is [^7] The results in this section hold for inference based on either Itô or Stratonovich calculus in [9]. Also, we have neglected the inference of the true mean $\langle\mathbf{x}\rangle$ and have assumed this as known to be $\mathbf{0}$. This simplifies the calculations while retaining interesting effects.:
@@
Perhaps surprisingly, there is no $d$-dependence to leading order in $T^{-1}$ (Fig. 3). This means that it is possible for the bias to dominate the variance in high dimensions. The $d$-dependence in the variance presumably occurs at higher order in $T^{-1}$, but we do not bother to calculate it.
@@ 3.2. Quantitative significance of the deterministic contribution
We see that depending on $D^{xy}$, the value can be either increased or decreased relative to that due to the deterministic ($\alpha$) or stochastic ($D^{yy}$) contributions alone. Nevertheless, some bounds can be established by using $\left|{D^{xy}}\right|\leq\sqrt{D^{xx}D^{yy}}$. By minimizing with respect to $\alpha$ while keeping $D^{yy}$ fixed, we get:
@@
Notice that both of these bounds stay finite as $\tau\to\infty$. The reason for this is that the $\tau\to\infty$ behavior of $\langle(y(\tau)-y(0))^{2}\rangle$ is given by $2D^{zz}\tau$ if $D^{zz}>0$, and $D^{zz}$ can be made arbitrarily small for any particular choice of either $\alpha$ and $D^{yy}$ (but not both simultaneously). Comparing these two bounds, we see that the relevant comparison is between $(\alpha/\lambda)\sqrt{D^{xx}}$ and $\sqrt{D^{yy}}$. Indeed, if $(|\alpha|/\lambda)\sqrt{D^{xx}}\ll\sqrt{D^{yy}}$, then the contribution of $D^{yy}$ to $\langle(y(\tau)-y(0))^{2}\rangle$ is dominant. On the other hand, if $(|\alpha|/\lambda)\sqrt{D^{xx}}\gg\sqrt{D^{yy}}$, then for long times $\tau\gg 1/\lambda$, the contribution of $\alpha$ to $\langle(y(\tau)-y(0))^{2}\rangle$ is dominant.
For multidimensional $\mathbf{x}$, we may compare the square roots of the quadratic variation of the two terms on the r.h.s. of Eq. (66).
@@ 3.3. Comparison of experimental data with theoretical predictions
To compare experimentally measured covariance functions to theoretical predictions, we can again normalize by the diffusion matrix. We may consider coordinates in which $\mathbf{D}=\mathds{1}$ and compare the experimental and theoretical values of $\langle\dot{\mathbf{x}}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle$. Or, for a coordinate-independent description, similarly to the case of stationary variables, we may decompose the deviation of experimental to theoretical $\langle\dot{\mathbf{x}}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle$ into symmetric and antisymmetric parts, multiply on the right by $\mathbf{D}^{-1}$ and inspect the eigenvalues and compare them to unity. In terms of “ensemble covariance”, this amounts to:
@@
where $\Delta_{\textrm{exp\textendash theo}}$ denotes the deviation of experimental to theoretical values. As mentioned for the angular momentum, if the time-step is not small, then on the r.h.s. of the above, it may be suitable to use discrete-time estimators of the diffusion. In case the eigenvalues of $\mathbf{H}$ exceed unity, the eigenvalues need to be compared to these instead, not just for the antisymmetric part but also for the symmetric part. We may illustrate this by way of two-dimensional linear Gaussian dynamics with complex conjugate eigenvalues $-\lambda\pm i\omega$ of $\mathbf{A}$ (Eq. (35)) and $D^{xx}=D^{yy}$, in which case:
@@
As an example, for a single (integrated) variable the quantity in Eq. (85) approaches as $\tau\to\infty$ the negative of the long-time diffusivity, to be compared to the short-time diffusivity. It should be noted, however, that the limit $\tau\to\infty$ cannot be taken for a trajectory of finite length. To illustrate this, suppose we have an integrated variable $x$ sampled with time-step $\Delta t$, and define $\Delta x_{n}:=x((n+1)\Delta t)-x(n\Delta t)$ for integer $n$, and $\Delta x_{n}^{\prime}:=\Delta x_{n}-\langle\Delta x\rangle$ (the subscript $n$ in the expectation value is suppressed). The long-time diffusivity is:
(86) $\lim_{N\to\infty}\frac{\langle(x(N\Delta t)-x(0)-\langle x(N\Delta t)-x(0)\rangle)^{2}\rangle}{2N\Delta t}$ $=\frac{1}{2N\Delta t}\lim_{N\to\infty}\sum_{0\leq m<N}\sum_{0\leq n<N}\langle\Delta x_{m}^{\prime}\Delta x_{n}^{\prime}\rangle$
@@ 4. Third-order properties
So far, we have dealt with linear Gaussian systems. These are characterized by the quantities $\mathbf{A}$ and $\mathbf{D}$ which give rise to the second-order moments $\langle\mathbf{x}\mathbf{x}^{\mathsf{T}}\rangle$ and $\langle\dot{\mathbf{x}}\mathbf{x}^{\mathsf{T}}\rangle$. To address the possibility of third-order moments, we may consider additional terms in the model:
@@
where all coefficients are constant, we have suppressed $t$-dependence on the r.h.s., and Einstein summation notation is used. Neither addition to the model can stand on its own: in general, $a^{i}_{jk}$ by itself leads to divergences of $\mathbf{x}\to\infty$, and $b^{ij}_{k}$ by itself renders $\langle\mathrm{d}[x^{i},x^{j}]/\mathrm{d}t\mid\mathbf{x}(t)\rangle$ not positive semidefinite. A third possible contribution,
@@
where $p(\mathbf{x})$ is the stationary distribution. Multiplying by an arbitrary function $f(\mathbf{x})$ and integrating, we get:
@@
For inhomogeneous diffusion, we have from Eq. (92):
@@ 5. Inhomogeneous diffusion
In this section, we shall restrict ourselves to the case where the drift $\bm{\mathcal{A}}(\mathbf{x})$ is a linear function $\mathbf{A}\mathbf{x}$. In addition to being easier to deal with, in the case of weak nonlinearity, coordinates can be chosen via Koopman eigenfunctions [19, 20] to transform the system into this form, as will be done in the next section. Additional terms in inhomogeneous diffusion are assumed to be small.
@@
Now we turn to an example in two dimensions. Assume that there is a symmetry $y\to-y$, so that $\mathbf{A}=\operatorname{diag}(-\lambda_{x},-\lambda_{y})$ with $D^{xy}=0$ and $b^{xx}_{y}=b^{xy}_{x}=b^{yy}_{y}=0$. Then $\langle x^{2}\rangle=D^{xx}/\lambda_{x}$, $\langle y^{2}\rangle=D^{yy}/\lambda_{y}$, and
@@
We see that there is only one “mode of time-reversal asymmetry”, corresponding to $L(x,y^{2})$. An angular momentum $L(x,y^{2})>0$ represents probability current which is anticlockwise for $y>0$ and clockwise for $y<0$. We see that $b^{yy}_{x}$ contributes positively to $L(x,y^{2})$. However, the sign of the contribution due to $b^{xy}_{y}$ depends on the ratio $\lambda_{x}/\lambda_{y}$.
@@
We can solve for the covariance functions in the same manner as before. Therefore, we focus on the condition for detailed balance. As before, we have:
@@
where we used $\widetilde{A}(\tau)^{ij}_{kl}=\widetilde{A}(\tau)^{ji}_{lk}$.
@@
If $\mathbf{A}\mathbf{C}=\mathbf{C}\mathbf{A}^{\mathsf{T}}$, it is seen that the two covariance functions are equal.
@@ 6.4. Generalized Koopman eigenfunctions
with $\langle\mathbf{x}(\tau)\mid\mathbf{x}(0)\rangle$ a quadratic function, since this would involve fifth moments.
@@ 6.5. Integrated variables
The above shows the values of $b^{iy^{\prime}}$. On the r.h.s., $x$ can be replaced by $x^{\prime}$ with an error $a\mathcal{O}(a,b)$. We can now calculate third-order covariance functions with $a\mathcal{O}((a,b)^{2})$ correction by evaluating fourth-order quantities by applying Isserlis’s theorem. The condition for detailed balance (with $a\mathcal{O}((a,b)^{2})$ correction) is also now ready to be formulated. When writing the time-reversibility for third-order covariance functions of the primed variables, one obtains those for the unprimed variables by adding $\mathcal{O}(a)$ terms involving fourth-order quantities. The fourth-order quantities satisfy time-reversibility with correction $\mathcal{O}((a,b)^{2})$ by applying Isserlis’s theorem and the time-reversibility of second-order moments. The condition for detailed balance then reduces to the time-reversibility for primed variables, which for the same reason reduces to the time-reversibility for unprimed variables.
@@ 6.7. Finite time interval inhomogeneous diffusion
However, for finite $\tau$, the above quantity can be non-zero for homogeneous diffusion because of non-vanishing third moments. This can be illustrated by the example of Eq. (90) in one dimension with $-A=D=1$. From Eqs. (153)–(154), we have:
@@ 6.8. Quantitative significance of third-order effects
Following the discussion of deviations of experiment from theory for the second moments, we may treat $(i,j,k)$ as a single “super-index” and similarly for $(i^{\prime},j^{\prime},k^{\prime})$, then multiply the l.h.s. by the matrix inverse of the r.h.s. For the same reason as before, the resulting matrix has a single non-zero eigenvalue $\mu$ equal to its trace. To satisfy symmetry, using the same argument as before, we may consider an ensemble of stochastic systems whose “ensemble covariance” of third moments (l.h.s. of the above) is:
@@
In the absence of symmetry constraints, we may calculate the “expected value” of $\mu$:
(191) $\mu=\langle x^{i}x^{j}x^{k}\rangle\langle x^{i^{\prime}}x^{j^{\prime}}x^{k^{\prime}}\rangle C^{-1}_{ii^{\prime}}C^{-1}_{jj^{\prime}}C^{-1}_{kk^{\prime}}\sim\frac{d^{3}+3d^{2}+2d}{6},$
@@
A putative relation for the “ensemble covariance” of the $b$ coefficients can be written as:
@@
Again, we may consider $(i,j)$ as a single “super-index” and similarly for $(i^{\prime},j^{\prime})$ and take the l.h.s. multiplied by the matrix inverse of the r.h.s. In this case, however, the resulting matrix will generally have multiple non-zero eigenvalues. Nevertheless, we may use the same arguments to deduce that the eigenvalues remain the same when symmetrizing the r.h.s. (which is inverted after neglecting permutations). In the absence of symmetry constraints, we may then consider the “ensemble covariance” for an ensemble of stochastic systems to be:
@@
where $d_{0}$ is the dimension of the stationary variables only and $d$ is the total dimension including integrated variables. In the above, $k$ and $k^{\prime}$ are summed over stationary variables only, whereas $i$, $j$, $i^{\prime}$, and $j^{\prime}$ are summed over all variables including integrated variables [^14] We may however be concerned about the possibility for a non-negative diffusivity function to have a mean far less than the standard deviation of its “linear component” (scaling as $\sqrt{d}$). To this end, we may consider the function $e^{cx}$, where $c$ is a constant, for $x\sim\mathcal{N}(0,1)$. Its mean value is $\langle e^{cx}\rangle=e^{c^{2}/2}$. When performing linear regression of $e^{cx}$ with regressor $x$, the coefficient is $\langle xe^{cx}\rangle=ce^{c^{2}/2}$. We see that the ratio between the two can be made arbitrarily large..
For the coefficients $a^{i}_{jk}$, we need to consider the matrix:
(195) $M^{jkj^{\prime}k^{\prime}}:=\langle(x^{j}x^{k}-C^{jk})(x^{j^{\prime}}x^{k^{\prime}}-C^{j^{\prime}k^{\prime}})\rangle,$
where $(j,k)$ is considered as a single “super-index”, as is $(j^{\prime},k^{\prime})$. We need to invert the above matrix neglecting permutations of $(j,k)$ and $(j^{\prime},k^{\prime})$. This is done by introducing a matrix $N_{jkj^{\prime}k^{\prime}}$ symmetric under the swappings $j\leftrightarrow k$, $j^{\prime}\leftrightarrow k^{\prime}$, and $(j,k)\leftrightarrow(j^{\prime},k^{\prime})$ such that:

### v14

@@ 2.4. Estimation of $\mathbf{A}$ from trajectories
We now consider the estimation of $\mathbf{A}$ in multiple dimensions. It has been shown that for estimation of the force field from stochastic trajectories, there is a finite rate at which information can be extracted [9]. This work has shown that the mean coefficient of the inferred force field $\widehat{\mathbf{A}}\mathbf{x}$ is $\langle\widehat{\mathbf{A}}\rangle=\mathbf{A}$ with correction $\mathcal{O}(T^{-1})$. However, they do not discuss the dependence of $\langle\widehat{\mathbf{A}}\rangle$ on dimension $d$. Eq. (C11) in [9] is summed over basis function index $\alpha$, and we may guess Eq. (C13) in [9] to be of order $\mathcal{O}(d^{3})$, leaving a $\mathcal{O}(d)$ bias for the inferred force field coefficient. To explore the impact of dimension, we explicitly calculate the bias and variance of $\widehat{\mathbf{A}}$ for a $d$-dimensional linear system in which $-\mathbf{A}=\mathbf{D}=\mathbf{C}=\mathds{1}$. According to the information-theoretic criterion in [9], the force field should start to be resolved for a trajectory length of 2 time-units. The value of $\widehat{\mathbf{A}}$ inferred from a trajectory of duration $T$ is [^8] The results in this section hold for inference based on either Itô or Stratonovich calculus in [9]. Also, we have neglected the inference of the true mean $\langle\mathbf{x}\rangle$ and have assumed this as known to be $\mathbf{0}$. This simplifies the calculations while retaining interesting effects.:
@@
Perhaps surprisingly, there is no $d$-dependence to leading order in $T$ (Fig. 3). This means that it is possible for the bias to dominate the variance in high dimensions. The $d$-dependence in the variance presumably occurs at higher order in $T^{-1}$, but we do not bother to calculate it.
@@ 3. Integrated variables
3.1. Preliminaries: Modeling, covariance functions, and detailed balance
@@ 3.2. Quantitative significance of the deterministic contribution
We see that depending on $D^{xy}$, the value can be either increased or decreased relative to that due to the deterministic ($\alpha$) or stochastic ($D^{yy}$) contributions alone. Nevertheless, some bounds can be established by using $\left|{D^{xy}}\right|\leq\sqrt{D^{xx}D^{yy}}$. We treat $\lambda$ and $D^{xx}$ as fixed. By minimizing with respect to $\alpha$ while keeping $D^{yy}$ fixed, we get:
@@
Notice that both of these bounds stay finite as $\tau\to\infty$. The reason for this is that the $\tau\to\infty$ behavior of $\langle(y(\tau)-y(0))^{2}\rangle$ is given by $2D^{zz}\tau$ if $D^{zz}>0$, and $D^{zz}$ can be made to be zero for any particular choice of either $\alpha$ and $D^{yy}$ (but not both simultaneously, because $D^{zz}\geq\left({\sqrt{D^{yy}}-|\alpha|\sqrt{D^{xx}}/\lambda}\right)^{2}$). Comparing these two bounds, we see that the relevant comparison is between $(\alpha/\lambda)\sqrt{D^{xx}}$ and $\sqrt{D^{yy}}$. Indeed, if $(|\alpha|/\lambda)\sqrt{D^{xx}}\ll\sqrt{D^{yy}}$, then the contribution of $D^{yy}$ to $\langle(y(\tau)-y(0))^{2}\rangle$ is dominant. On the other hand, if $(|\alpha|/\lambda)\sqrt{D^{xx}}\gg\sqrt{D^{yy}}$, then for long times $\tau\gg 1/\lambda$, the contribution of $\alpha$ to $\langle(y(\tau)-y(0))^{2}\rangle$ is dominant.
For multidimensional $\mathbf{x}$, we may compare the square roots of the quadratic variation of the two terms on the r.h.s. of Eq. (66), i.e., $\sqrt{D^{yy}}$ with $\sqrt{\bm{\alpha}^{\mathsf{T}}\mathbf{A}^{-1}\mathbf{D}^{\mathbf{xx}}(\mathbf{A}^{-1})^{\mathsf{T}}\bm{\alpha}}$.
@@ 3.3. Comparison of experimental data with theoretical predictions
To compare experimentally measured covariance functions to theoretical predictions, we can again normalize by the diffusion matrix. We may consider coordinates in which $\mathbf{D}=\mathds{1}$ and compare the experimental and theoretical values of $\langle\dot{\mathbf{x}}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle$. In terms of the “ensemble covariance”, this amounts to:
@@
where $\Delta_{\textrm{exp\textendash theo}}$ denotes the deviation of experimental to theoretical values. As mentioned for the angular momentum, if the time-step is not small, then on the r.h.s. of the above, it may be suitable to use discrete-time estimators of the diffusion. In case the angular momenta exceed the “reference” values (Eq. (43)) by some factor, we may need to relax the above comparison by the same factor, not just for the antisymmetric part but also for the symmetric part. We may illustrate this by way of two-dimensional linear Gaussian dynamics with complex conjugate eigenvalues $-\lambda\pm i\omega$ of $\mathbf{A}$ (Eq. (34)) and $D^{xx}=D^{yy}$, in which case:
@@
Thus if $\omega/\lambda\gtrsim 1$, deviation of experimental to theoretical values could be amplified by this factor without necessarily being considered quantitatively significant. Similarly to the comparison of experimental values to theoretical predictions of $\langle\mathbf{x}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle$, we may compare separately the symmetric and antisymmetric components (with respect to indices), which correspond to time-symmetric and -antisymmetric quantities, respectively.
@@
As an example, for a single (integrated) variable the quantity in Eq. (85) approaches as $\tau\to\infty$ the negative of the long-time diffusivity, to be compared to the short-time diffusivity.
It should be noted, however, that the limit $\tau\to\infty$ cannot be taken for a trajectory of finite length. To illustrate this, suppose we have an integrated variable $x$ sampled with time-step $\Delta t$, and define $\Delta x_{n}:=x((n+1)\Delta t)-x(n\Delta t)$ for integer $n$, and $\Delta x_{n}^{\prime}:=\Delta x_{n}-\langle\Delta x\rangle$ (the subscript $n$ in the expectation value is suppressed). The long-time diffusivity is:
(86) $\lim_{N\to\infty}\frac{\langle(x(N\Delta t)-x(0)-\langle x(N\Delta t)-x(0)\rangle)^{2}\rangle}{2N\Delta t}$ $=\lim_{N\to\infty}\frac{1}{2N\Delta t}\sum_{0\leq m<N}\sum_{0\leq n<N}\langle\Delta x_{m}^{\prime}\Delta x_{n}^{\prime}\rangle$
@@ 4. Third-order properties
So far, we have dealt with linear Gaussian systems. These are characterized by the quantities $\mathbf{A}$ and $\mathbf{D}$ which give rise to the second-order moments $\langle\mathbf{x}\mathbf{x}^{\mathsf{T}}\rangle$ and $\langle\dot{\mathbf{x}}\mathbf{x}^{\mathsf{T}}\rangle$. In this section, we will establish general properties of Langevin equations beyond second order. To address the possibility of third-order effects, we may consider additional terms in the model:
@@
where all coefficients are constant, we have suppressed $t$-dependence on the r.h.s., and Einstein summation notation is used. Neither addition to the model can stand on its own: in general, $a^{i}_{jk}$ by itself leads to divergences of $\mathbf{x}\to\infty$, and $b^{ij}_{k}$ by itself renders $\langle\mathrm{d}[x^{i},x^{j}]/\mathrm{d}t\mid\mathbf{x}(t)\rangle$ not positive semidefinite. In section 6, we will discuss the magnitude of additional terms needed to result in an admissible model. A third possible contribution,
@@
where $p(\mathbf{x})$ is the stationary distribution and $\partial_{j}:=\partial/\partial x^{j}$. Multiplying by an arbitrary function $f(\mathbf{x})$ and integrating, we get:
@@
Thus $L(x^{i}x^{j},y)$ are the only third-order quantities odd under time reversal involving an integrated variable. For inhomogeneous diffusion, we have from Eq. (92):
@@
Thus in the expression $\langle x^{i}\mathrm{d}[x^{j},x^{k}]/\mathrm{d}t\rangle$, the time at which $x^{i}$ is evaluated has no impact.
@@ 5. Inhomogeneous diffusion
In this section, we shall restrict ourselves to the case where the drift $\bm{\mathcal{A}}(\mathbf{x})$ is a linear function $\mathbf{A}\mathbf{x}$. In addition to being easier to deal with, in the case of weak nonlinearity, coordinates can be chosen via Koopman eigenfunctions [19, 20] to transform the system into this form, as will be done in the next section. Additional terms in inhomogeneous diffusion are assumed to be small. In this section, we will discuss the effect of inhomogeneous diffusion on third-order covariance functions.
5.1. Stationary case
@@
5.2. Two-dimensional example
Now we turn to an example in two dimensions, with variables $x$ and $y$. Assume that there is a symmetry $y\to-y$, so that $\mathbf{A}=\operatorname{diag}(-\lambda_{x},-\lambda_{y})$ with $D^{xy}=0$ and $b^{xx}_{y}=b^{xy}_{x}=b^{yy}_{y}=0$. Then $\langle x^{2}\rangle=D^{xx}/\lambda_{x}$, $\langle y^{2}\rangle=D^{yy}/\lambda_{y}$, and
@@ 5.2. Two-dimensional example
We see that there is only one “mode of time-reversal asymmetry”, corresponding to $L(x,y^{2})$. An angular momentum $L(x,y^{2})>0$ represents probability current which is anticlockwise for $y>0$ and clockwise for $y<0$. We see that $b^{yy}_{x}$ and $b^{xy}_{y}$ contribute positively to $\langle xy^{2}\rangle$, and that $b^{yy}_{x}$ contributes positively to $L(x,y^{2})$. However, the sign of the contribution of $b^{xy}_{y}$ to $L(x,y^{2})$ depends on the ratio $\lambda_{x}/\lambda_{y}$.
@@ 5.3. Integrated variables
When stationary variables are expressed in “covariance-identity” coordinates, $b^{iy}_{j}$ and $b^{jy}_{i}$ both contribute positively to $L(x^{i}x^{j},y)$.
We can solve for the covariance functions in the same manner as before. Therefore, we focus on the condition for detailed balance. Similarly to before, we have:
@@
 [^14] The limit is needed because $\mathbf{A}\otimes\mathds{1}-\mathds{1}\otimes\mathbf{A}$ is not invertible. To see this, let $\mathbf{v}$ be an eigenvector of $\mathbf{A}$. Then $(\mathbf{A}\otimes\mathds{1}-\mathds{1}\otimes\mathbf{A})(\mathbf{v}\otimes\mathbf{v})=\mathbf{0}$.
@@
where we used $\widetilde{A}(\tau)^{ij}_{kl}=\widetilde{A}(\tau)^{ji}_{lk}$. Thus we see that under assumption of detailed balance to second and third order, the investigated third-order covariance functions are time-reversible.
@@
If also $\mathbf{A}\mathbf{C}=\mathbf{C}\mathbf{A}^{\mathsf{T}}$, it is seen that the two covariance functions are equal.
@@ 6.1. Preliminaries
The multidimensional version is given by Eqs. (90)–(91). In sections 6.2–6.6, we will present the solution of this model, and in sections 6.7–6.10, we will discuss inference, quantitative significance, and comparison between theory and experiment.
@@ 6.4. Generalized Koopman eigenfunctions
with $\langle\mathbf{x}(\tau)\mid\mathbf{x}(0)\rangle$ a quadratic function, since this would involve fifth moments. Also, note that if $\mathbf{A}$ has eigenvalues $\lambda_{1},\lambda_{2},\lambda_{3}$ such that $\lambda_{1}=\lambda_{2}+\lambda_{3}$, then $\mathbf{A}\otimes\mathds{1}\otimes\mathds{1}-\mathds{1}\otimes\mathbf{A}\otimes\mathds{1}-\mathds{1}\otimes\mathds{1}\otimes\mathbf{A}$ is not invertible [^16] If $\mathbf{v}_{i}$ ($i=1,2,3$) are eigenvectors of $\mathbf{A}$ with eigenvalues $\lambda_{i}$, respectively, then $(\mathbf{A}\otimes\mathds{1}\otimes\mathds{1}-\mathds{1}\otimes\mathbf{A}\otimes\mathds{1}-\mathds{1}\otimes\mathds{1}\otimes\mathbf{A})(\mathbf{v}_{1}\otimes\mathbf{v}_{2}\otimes\mathbf{v}_{3})=\mathbf{0}$. and Koopman eigenfunctions cannot in general be found.
@@ 6.5. Integrated variables
The above shows the values of $b^{iy^{\prime}}$. On the r.h.s., $x$ can be replaced by $x^{\prime}$ with an error $a\mathcal{O}(a,b)$. We can now calculate third-order covariance functions with $a\mathcal{O}((a,b)^{2})$ correction by evaluating fourth-order quantities by applying Isserlis’s theorem. The condition for detailed balance (with $a\mathcal{O}((a,b)^{2})$ correction) is also now ready to be formulated. When writing the time-reversibility for third-order covariance functions of the primed variables, one obtains those for the unprimed variables by adding $\mathcal{O}(a)$ terms involving fourth-order quantities. The fourth-order quantities satisfy time-reversibility with correction $\mathcal{O}((a,b)^{2})$ by applying Isserlis’s theorem and the time-reversibility of second-order moments. The condition for detailed balance then reduces to the vanishing of the angular momenta for primed variables, which for the same reason reduces to the vanishing of the angular momenta for unprimed variables.
@@ 6.7. Finite time interval inhomogeneous diffusion
However, for finite $\tau$, the above quantity can be non-zero for homogeneous diffusion because of non-vanishing third moments. This can be illustrated by the example of Eq. (90) in one dimension with $-A=D=1$ and $b=0$. From Eqs. (153)–(154), we have:
@@ 6.8. Quantitative significance of third-order effects
Following the discussion of deviations of experiment from theory for the second moments, we may treat $(i,j,k)$ as a single “super-index” and similarly for $(i^{\prime},j^{\prime},k^{\prime})$, then multiply the l.h.s. by the matrix inverse of the r.h.s. For the same reason as before, the resulting matrix has at most one non-zero eigenvalue equal to its trace. To satisfy symmetry, using the same argument as before, we may consider an ensemble of stochastic systems whose “ensemble covariance” of third moments (l.h.s. of the above) is:
@@
In the absence of symmetry constraints, we may calculate the “expected value”:
(191) $\langle x^{i}x^{j}x^{k}\rangle\langle x^{i^{\prime}}x^{j^{\prime}}x^{k^{\prime}}\rangle C^{-1}_{ii^{\prime}}C^{-1}_{jj^{\prime}}C^{-1}_{kk^{\prime}}\sim\frac{d^{3}+3d^{2}+2d}{6},$
@@
Now, we turn to the coefficients for inhomogeneous diffusion. A putative relation for the “ensemble covariance” of the $b$ coefficients can be written as:
@@
For fixed $k$ and $k^{\prime}$, we may consider $(i,j)$ as a single “super-index” and similarly for $(i^{\prime},j^{\prime})$ and take the l.h.s. multiplied by the matrix inverse of the r.h.s. Similarly to before, in coordinates where $\mathbf{D}=\mathds{1}$ (without transforming $k$ and $k^{\prime}$), the resulting matrix has the form $\mathbf{v}\mathbf{w}^{\mathsf{T}}$ and therefore has rank at most 1. Thus, we may use the same argument to deduce that the eigenvalues not associated with antisymmetric left eigenvectors remain the same when symmetrizing the r.h.s., which is inverted after neglecting permutations. In the absence of symmetry constraints, we may then consider the “ensemble covariance” for an ensemble of stochastic systems to be:
@@
where $d_{0}$ is the dimension of the stationary variables only and $d$ is the total dimension including integrated variables. In the above, $k$ and $k^{\prime}$ are summed over stationary variables only, whereas $i$, $j$, $i^{\prime}$, and $j^{\prime}$ are summed over all variables including integrated variables [^17] We may however be concerned about the possibility for a non-negative diffusivity function to have a mean far less than the standard deviation of its “linear component” (scaling as $\sqrt{d}$). To this end, we may consider the function $e^{cx}$, where $c$ is a constant, for $x\sim\mathcal{N}(0,1)$. Its mean value is $\langle e^{cx}\rangle=e^{c^{2}/2}$. When performing linear regression of $e^{cx}$ with regressor $x$, the coefficient is $\langle xe^{cx}\rangle=ce^{c^{2}/2}$. We see that the ratio between the two can be made arbitrarily large, and thus there is no issue.. The r.h.s. is the number of degrees of freedom of $b^{ij}_{k}$. As for second-order quantities, if the time-step is not small, it may be suitable to use discrete-time estimators of the diffusion.
Next, we address the nonlinear drift. For the coefficients $a^{i}_{jk}$, we would like to compare $a^{i}_{jk}(x^{j}x^{k}-C^{jk})$ with $A^{i}_{j}x^{j}$. We are thus led to consider the matrix:
(195) $M^{jkj^{\prime}k^{\prime}}:=\langle(x^{j}x^{k}-C^{jk})(x^{j^{\prime}}x^{k^{\prime}}-C^{j^{\prime}k^{\prime}})\rangle=C^{jj^{\prime}}C^{kk^{\prime}}+C^{jk^{\prime}}C^{kj^{\prime}},$
evaluated in the linear Gaussian model, where $(j,k)$ is considered as a single “super-index”, as is $(j^{\prime},k^{\prime})$. We need to invert the above matrix neglecting permutations of $(j,k)$ and $(j^{\prime},k^{\prime})$. This is done by introducing a matrix $N_{jkj^{\prime}k^{\prime}}$ symmetric under the swappings $j\leftrightarrow k$, $j^{\prime}\leftrightarrow k^{\prime}$, and $(j,k)\leftrightarrow(j^{\prime},k^{\prime})$ such that: